# Gaussian Splatting in Style

### v1

@@ 5 Experiments
Dataset
Method
GS →
AdaIN
AdaIN
→ GS
LSNV ARF
Stylised-
NeRF
StyleRF GSS(Ours)
RMSE ↓ 0.098 0.083 0.030 0.036 0.002 0.038 0.034
Flower LPIPS ↓ 0.106 0.080 0.011 0.015 0.001 0.013 0.009
RMSE ↓ 0.127 0.116 0.075 0.038 0.041 0.037 0.026
Fern LPIPS ↓ 0.151 0.134 0.049 0.020 0.012 0.019 0.008
RMSE ↓ 0.104 0.103 0.044 0.037 0.029 0.035 0.021
Room LPIPS ↓ 0.109 0.105 0.014 0.014 0.011 0.014 0.006
RMSE ↓ 0.136 0.099 0.089 0.043 0.015 0.045 0.010
Trex LPIPS ↓ 0.136 0.087 0.033 0.022 0.004 0.017 0.003
Table 1: Short Term Consistency Here, we demonstrate the superior performance of our method using the short-term consistency metric. We compute the metric using both, RMSE and LPIPS[88] using a VGG backbone. As shown in Figure 4, the novel views rendered by StylizedNerf are highly blurry and misses crucial details. Therefore, due to the over-smooth and blurry output, the metric obtained for them is really low and does not correctly reflect the adequate stylization qualities.
Implementation Details We build our work on top of the 3DGS[33] framework and use it to pretrain the Gaussians for each scene. For this, we followed the original training regime as suggested by the authors. For the 2D stylization module, we use a pretrained VGG [72, 16] and AdaIN[27] decoder. For obtaining the latent code of input style images, we use a pretrained image encoder used in [65]. As for the hyperparameters, we set $\lambda_{g}$ and $\lambda_{c}$ as 1 and 0.1 respectively. We train the tiny MLP for the second stage as per the training regime described in Figure 3 for 50k iterations. For optimization, we use the Adam [34] optimizer with a learning rate of $1e^{-4}$. We run all our experiments on a Nvidia A4500 GPU.
Baselines Since no previous approach performs scene stylization based on Gaussian splatting, we compare our method with NeRF-based approaches. We primarily compare against StylizedNerf[28, 24], and LSNV[26] as their problem statement resembles most closely to that of ours. We also compare against a newer method StyleRF[43], along with artistic radiance fields(ARF)[87], although ARF optimizes every scene for each given style image. For completeness, we also create two other baselines, ADA-GS and GS-ADA. In ADA-GS, we train 3DGS on 2D stylized images for each style. This setting resembles that of ARF. In GS-ADA, we apply AdaIN[27] on the novel view renders obtained from pretrained 3DGS.
@@
Datasets Similar to the baselines [28, 43, 26] we perform stylization on various real-world datasets. For real-world, we use the LLFF dataset[49]. It contains high-resolution front-facing scenes, out of which we select the most common ones used in the baselines, namely $trex,fern,flowers,room$. We also demonstrate the effectiveness of our method on more challenging Tanks and Temples(TnT)[36] dataset. For training, we use a dataset of 120 diverse style images sampled from WikiArt dataset[68], which include the paintings and popular artworks of verious artists.
Dataset
Method
GS →
AdaIN
AdaIN
→ GS
LSNV ARF
Stylised-
NeRF
StyleRF GSS(Ours)
RMSE ↓ 0.147 0.141 0.040 0.123 0.002 0.123 0.090
Flower LPIPS ↓ 0.151 0.135 0.018 0.112 0.001 0.081 0.057
RMSE ↓ 0.151 0.132 0.076 0.109 0.044 0.115 0.042
Fern LPIPS ↓ 0.193 0.162 0.051 0.107 0.019 0.099 0.021
RMSE ↓ 0.143 0.150 0.045 0.084 0.045 0.078 0.033
Room LPIPS ↓ 0.182 0.184 0.019 0.061 0.028 0.046 0.015
RMSE ↓ 0.164 0.131 0.091 0.092 0.030 0.091 0.020
Trex LPIPS ↓ 0.194 0.137 0.050 0.081 0.016 0.057 0.007
Table 2: Long Term Consistency In this table, we measure the performance of GSS against the said baselines via long-term consistency. Similar to [43], for long-term consistency, the two frames should be far away from each other. Having a long-term consistent method ensures that the produced novel views of the scene are consistent in the stylized color space. Similar to short-term consistency in Table 1, StylizedNerf numbers does not accurately depict the stylization quality as compared to other methods.
5.1 Quantitative Results
@@
Since the task of scene stylization is relatively new, there are not many ways to quantitatively compare methods. However, following [26, 28, 43], we measure the quantitative result by measure short-term and long-term consistency. Similar to the evaluation in [43], we first take two stylized views. By using [76] and [58], we then warp one view into another before computing the masked LPIPS[88] and RMSE score. We use adjacent views for short-term and far-away views for long-term consistency. Many works such as [26] also refer to it as warped LPIPS and warped RMSE. Mathematically, they can be summarised as follows:
@@
$\mathbb{E}_{wrmse}(\mathcal{O}_{v},\mathcal{O}_{v^{{}^{\prime}}})=RMSE(\mathcal{O}_{v},\mathcal{M}_{v}(\mathcal{W}(\mathcal{O}_{v^{{}^{\prime}}})))$ (8)
where $\mathcal{M}_{v}$, $\mathcal{O}_{v}$, and $\mathcal{W}$ are the masking, view ${v}$, and the warping function respectively.
We quantitatively compare our method with the baselines as shown in Tables 1 and 2. It can be clearly seen that our method outperforms all other baselines in both short-term and long-term consistency. We can see that simply applying 2D stylization to rendered novel views is not enough to ensure consistency, in accordance with [26], and hence Ada-GS is almost always more consistent than GS-ADA for both short-term and long-term views. Furthermore, simply adding 2D stylization to the input or outputs of existing scene representation approaches is not sufficient, as all other dedicated baselines perform better on average than GS-ADA and ADA-GS. We also find that methods utilising explicit scene representations, such as GSS and LSNV[26] perform better than radiance field based baselines. It is also important to observe that for StylizedNerf, the generated novel views suffer from high blur, over-smoothing and loss of details. Due to this, they obtain a very high consistency despite not being of reasonable quality as compared to all other methods.
@@
On running on the same hardware, GSS is also the fastest. Whereas all other methods take in the order of a few seconds to minutes to generate a single view, we obtain a rendering speed of roughly averaging 157 FPS. Our method is also very efficient in training and it requires roughly 2 hours to train for each scene on a standard Nvidia A4500 GPU.
Figure 4: We provide a detailed qualitative comparison of our method against the baselines detailed in Section 5. It can be seen that GSS achieves a highly accurate stylization based on the input style. While methods such as ARF obtain a better texture, it is attributed to the fact that it is optimized separately for each style. GS-Ada and Ada-GS are also able to get high quality texture, however, it is not accurate at geometry conservation since these methods are based on pure 2D stylization. Also, methods such as StyleRF fail to grasp the accurate style color such as that in room, fern or flower scene.
5.2 Qualitative Results
@@ 6 Ablation Studies
Figure 5: Here we show the effect of varying $\lambda_{c}$ in our loss function as described in Equation 5. We observe that setting $\lambda_{c}=0$ results in loss of details such as the anthers and filaments of the flower as shown above. However, further increasing $\lambda_{c}$ results in loss of style, and the generated view looks similar to that of the unstylized view.
6.1 Content Loss
We first ablate the content loss $\mathcal{L}_{c}$, described in Equation 5, by varying the value of $\lambda_{c}$. We vary the value in the discrete set of $\{0,0.1,1,10,100\}$ to cover a wide range of possible values.
Setting the value of $\lambda_{c}=0$ leads to over-smoothing of the generated view. While the generated view is still consistent and stylized, it loses some of the finer details. On increasing the value beyond 1, the rendered view resembles that of being unstylized.
Figure 6: We ablate the effect of multi-resolution hash grid(MHG) in our setup. We find that having MHG in our model enforces a much better texture and color transfer from the style image onto the scene. It can be attributed to its high capability of learning features, while keeping the model fast and small.
6.2 Multi-resolution Hash Grid
We also study the effect of having a multi-resolution hash grid(MHG). On removing the MHG from our 3D color module, we find that the resulting stylized scene is not adequately transferred onto the scene as shown in Section 6.1. Instead, the scene loosely adopts the color-palette from the style image without grasping any textual information, thus verifying our decision to use MHG.

### v2

@@ 5 Experiments
Table 1: Here, we demonstrate the superior performance of our method using the short-term and long-term consistency metric. We compute the metric using a VGG backbone for both, RMSE and LPIPS[88], elaborated in Section 5.2.1 . We provide a qualitative evaluation to support visual superiority of our method against the mentioned baselines in Figure 4, and also provide videos for a comprehensive evaluation in the supplementary materials.
LLFF[49] TnT[36]
Short-Term Long-Term Short-Term Long-Term
Method RMSE LPIPS RMSE LPIPS RMSE LPIPS RMSE LPIPS
GS $\rightarrow$ AdaIN 0.122 0.132 0.153 0.190 0.050 0.014 0.107 0.054
AdaIN $\rightarrow$ GS 0.106 0.109 0.138 0.161 0.015 0.002 0.100 0.055
LSNV [26] 0.069 0.032 0.070 0.040 0.030 0.006 0.100 0.045
ARF [87] 0.039 0.019 0.095 0.083 0.012 0.001 0.101 0.048
Stylised-NeRF [28] 0.028 0.009 0.040 0.021
StyleRF [43] 0.039 0.017 0.095 0.067 0.024 0.006 0.130 0.070
Ours 0.019 0.005 0.032 0.014 0.015 0.002 0.083 0.040
In this section, we first present the implementation details of our method, followed by comparisons with existing methods and ablation studies.
5.1 Implementation Details
Our implementation is based on the 3DGS [33] framework. We closely follow the implementation of the method. We employ a warmup strategy, similar to Lee et al. [38], wherein we train 3DGS without our 2D stylization and the 3D Color modules. This is so that consistency is established in the initial iterations of the training and initialized Gaussians. We do this for the first 3000 iterations. Similar to 3DGS training, we allow the Gaussians to densify until 15000 iterations, after which we fix the number of Gaussians in the scene. For the 2D stylization module, we use a pre-trained VGG model [72, 16] and AdaIN decoder [27]. We encode the input style images using a pre-trained image encoder  [65]. We train an MLP after the warmup as described in “training regime” for a total of 150k iterations. The MLP is composed of 2 fully connected layers with 64 neurons each. For the MHG, we use a resolution of 16 levels, with each level having a feature vector of length 2. For optimization, we use the Adam optimizer  [34] with a learning rate of $1e^{-4}$. We run all our experiments on a Nvidia A4500 GPU.
5.2 3D Stylization with Gaussian Splatting
We compare our method on stylizing 3D scenes against StylizedNerf[28, 24], LSNV[26], StyleRF[43], and artistic radiance fields(ARF)[87]. We would like to emphasize that StylizedNerf, StyleRF, and LSNV are more relevant to our method as the methods focus on generalizing to multiple styles. ARF is optimized on a given style per scene. Further, we compare with two additional baselines that we call Ada-GS and GS-Ada. In Ada-GS, we train 3DGS on 2D stylized images for each style. This setting resembles that of ARF. In GS-Ada, we apply AdaIN[27] on novel views rendered from a pre-trained 3DGS model.
@@ 5.1 Implementation Details
Datasets Following existing methods [28, 43, 26], we stylize various real-world datasets, LLFF dataset [49], and Tanks and Temples (TnT) [36]. LLFF dataset consists of high-resolution forward-facing images of 3D scenes. We stylize trex, fern, and room from the LLFF dataset. We learn to stylize each scene over a dataset of approximately 20000 diverse style images sampled from the WikiArt dataset [68] consisting of paintings and popular artworks of various artists.
5.2.1 Quantitative Results
@@
2D or 3D Stylization, in general, is a subjective task. Hence, no metrics exist that can effectively measure stylization. However, we our goal is to stylize images in a 3D consistent way. Therefore, similar to existing works [26, 28, 43], we measure short-term and long-term consistency. We use adjacent views for short-term and far-away views for long-term consistency. Similar to the evaluation in [43], we first take two stylized views. By using optical flow based methods[76, 58], we then warp one view into another before computing the masked LPIPS[88] and RMSE score. Many works such as [26] also refer to it as warped LPIPS and warped RMSE. Mathematically, they can be summarised as
@@
and
$\mathbb{E}_{wrmse}(\mathcal{O}_{v},\mathcal{O}_{v^{{}^{\prime}}})=RMSE(\mathcal{O}_{v},\mathcal{M}_{v}(\mathcal{W}(\mathcal{O}_{v^{{}^{\prime}}}))),$ (6)
where $\mathcal{M}_{v}$, $\mathcal{O}_{v}$, and $\mathcal{W}$ are the masking, the rendered view, and the warping function respectively for two views ${v}$ and ${v}^{{}^{\prime}}$.
We quantitatively compare our method with the baselines as shown in Table 1. It can be clearly seen that our method outperforms all other baselines in both short-term and long-term consistency. We can see that simply applying 2D stylization to rendered novel views is not enough to ensure consistency, in accordance with [26], and hence Ada-GS is almost always more consistent than GS-ADA for both short-term and long-term views. Furthermore, simply adding 2D stylization to the input or outputs of existing scene representation approaches is not sufficient, as all other dedicated baselines perform better on average than GS-ADA and ADA-GS. We also find that methods utilising explicit scene representations, such as GSS and LSNV[26] perform better than radiance field based baselines. It is also important to observe that for StylizedNerf, the generated novel views suffer from high blur, over-smoothing and loss of details. Due to this, they obtain a very high consistency despite not being of reasonable quality as compared to all other methods.
@@
On running on the same hardware, GSS is also the fastest. Whereas all other methods take in the order of a few seconds to minutes to generate a single view, we obtain a rendering speed of roughly averaging 157 FPS. Our method is also efficient in training and it requires roughly 4 hours to train for each scene on a standard Nvidia A4500 GPU.
[…]
Figure 4: We provide a detailed qualitative comparison of our method against the baselines detailed in Section 5 on the (a) TnT and (b) LLFF datasets. Our method achieves a highly accurate stylization based on the input style. While methods such as ARF obtain a better texture, it is attributed to the fact that it is optimized separately for each style. StylizedNerf[28] produces images that suffer from over smoothing and blurriness, while StyleRF fails to grasp the accurate style color. On the other hand, our proposed method is able to retain high details present in the unstyled view while transferring the adequate texture and colors of the style image for both, indoor and outdoor real-world datasets.
5.2.2 Qualitative Results
@@ 6 Ablation Studies
6.0.1 Effect of Joint Training
[…]
Figure 5: We show the effect of deploying a joint-training regime of training the 3D Gaussians in conjunction with the 3D Color module. Having a joint training in an end-to-end fashion helps to preserve key details and geometry in the rendered stylized novel view.
Contrary to previous NeRF-based baselines such as StylizedNerf[28], we employ a joint-training regime where the 3D Gaussians of each scene are trained in conjuction with the 3D color module. To study the effect of this regime, we perform an ablation where we employ pre-trained 3D Gaussians. These Gaussians were trained on unstylized images, and were fixed throughout the training process while the 3D Color module was trained from scratch. The stylization results from pretrained Gaussians appear blurry while our method retains significantly higher amount of details. Furthermore, since the geometric parameters such as opacity are fixed in the pretrained model, the final style imparted onto the rendered view lacks coherence.
6.0.2 Train on stylized images vs stylizing novel renderings
Additionally, we quantitatively and qualitatively evaluate the performance of our synthetic baselines, namely GS-Ada and Ada-GS in Figures 6 and 1. We elaborate on these two baselines in Section 5.2. We observe that GS-Ada is the least consistent as the rendered novel views are stylized independently of each other using AdaIN [27], therefore, have no notion of consistency. While Ada-GS performs better, it has a major drawback, i.e. the method needs to be trained from scratch for every new style image, thus limiting its usability, especially for practical applications.
[…]
Figure 6: In addition to providing quantitative results in Table 1, we provide the qualitative results showing the distored geometry when we apply 2D stylization on rendered novel views(GS-Ada). We also observe that training vanilla 3DGS on 2D Stylized views(Ada-GS) preserves better details than GS-Ada. However, our method provides the most geometrically accurate renderings while being truthful to the queried style.